Stone Sour Co. has an ROA of9 percent and a payout ratio of 30 percent hat is its internal growth rate? (Do not round intermediate calculations and enter your answer as a percent Tounded to 2 decimal places:e0 92.16 Internal growth rate
Answer #1

Retention ratio=1-payout ratio

=(1-0.3)=0.7

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

=(0.09*0.7)/[1-(0.09*0.7)]

which is equal to

=6.72%(Approx).